How long will it take for prices in the economy to double at a 6% annual inflation rate? Round to the nearest hundredth.
Answer by Theo(13342)

F = P * (1.06)^x

Let P = 1
Let F = 2 * 1

Formula becomes:

2 = 1 * (1.06)^x

This is the same as:

2 = (1.06)^x

Take the log of both sides to get:

log(2) = log(1.06^x)

This is the same as:

log(2) = x*log(1.06)

divide both sides of this equation by log (1.06) to get:

log(2) / log(1.06) = x

Use log function of your calculator to solve to get:

x = 11.89566105

Confirm by substituting in original equation to get:

2 = (1.06)^(11.89566105) which becomes:

2 = 2 confirming that x = 11.89566105 is good.

your answer is 11.90 years.